mhchen:

Vector space: Let a and b be vectors. Define addition a + b = (a1,a2) + (b1,b2) = (a1+b1+1,a2+b2+1) and k(a) = (k*a1,k*a2) Show that (-1,-1) is the zero vector.

mhchen:

The definition of zero vector (provided by my textbook) is that 0 + 0 = 0 and k0 = 0 for all scalars k. It is easy to show how 0 + 0 = 0 But assuming (-1,-1) is the zero vector, then k0 = k(-1,-1) = (-k,-k) which does not equal the 0 vector?? Also I must show that a + (-a) = (-a) + a = 0 But a + (-a) = (1,1)

Gdeinward:

AHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHH!!!!!!. A.... A.... A PROOOOOOOOFFF